OpenStudy (anonymous):

Which equation has no solution? (1 point) 5v + 3 = 5v – 5 2x – 4 = 2x + 7 – x 2y + 5 – y = 2y – 3(y – 6) 7z + 9 = –7z – 6

OpenStudy (rushwr):

the 1st one

OpenStudy (texaschic101):

yep ^^

OpenStudy (rushwr):

5v +3 = 5v-5 +3 = 5v- 5v -5 3 = 5 is the answer u get which is obviously not the answer !

OpenStudy (anonymous):

Solve the following equation. 2(h – 8) – h = h – 16 (1 point) h = 8 h = –8 infinitely many solutions no solution

OpenStudy (texaschic101):

2(h - 8) - h = h - 16 --- distribute through the parenthesis 2h - 16 - h = h - 16 -- simplify h - 16 = h - 16 whenever your equations ends like this....where both sides are equal to each other, this means that the equation has infinite solutions. No matter what number you sub in for x, the equation will be true.

OpenStudy (anonymous):

Solve the following equation. 4 + 2y = 9 + 2y (1 point) 4 –4 infinitely many solutions no solution